OpenStudy (anonymous):

Did I do this correct??? Slope 2/5, through (0,-1) Y-Y1 = M (X-X1) Y-(-1) = 2/5 [x-(0)] Y - (-1) = 2/5x - 0 Y= 2/5x + 1

OpenStudy (anonymous):

Y - (-1) = 2/5x - 0 y+1=(2/5)x y=(2/5)x -1

OpenStudy (anonymous):

If the questions asks "Find the equation of the line with the given properties" (which is what I just did in the above problem) but the word states "contains (-2,6) and (-1,8) is it asking me the same thing?